# Can AI writing be salvaged? Mitigating Idiosyncrasies and Improving Human-AI Alignment in the Writing Process through Edits

### v3

@@ 7.2. How well can LLMs mimic edits from writers?
Our preference ranking results in Section 6.3 indicate that automatically edited paragraphs frequently rank second and sometimes first. This raises questions about LLMs’ ability to analyze textual patterns and generate content closely resembling a given writer’s edit. For the span “Janet lay in bed each night, her mind a whirlpool of restless thoughts”, both LLM and writer identified it as cliché. The LLM edited it to “Janet lay in bed each night, unable to sleep” while the writer changed it to “Each night, Janet lay prone in her bed and unable to sleep”. LLMs can also split run-on sentences and improve poor structure. For “Sarah froze, realizing it was her high school sweetheart, Alex, whom she hadn’t seen in over a decade”, the LLM edited it to “Sarah froze. It was Alex, her high school sweetheart. She hadn’t seen him in over a decade”, similar to the writer’s edit. However, LLMs sometimes replace clichés with other clichés or fail to remove unnecessary exposition. The most challenging edit category is Lack of Specificity and Detail where LLMs often fail to add engaging details. For “Her irritation slowly morphed into a strange, disconnected calm”, the writer added “After all, the noise just meant that she wasn’t the only one awake at this hour.” The model’s edit was less effective: “Her irritation slowly morphed into a strange, disconnected calm. The repetitive thump-thump-thump became almost hypnotic, lulling her into a trance-like state”. One potential limitation in our experiments is our reliance on few-shot instructions, requiring the model to learn rewriting from only a few examples. Training on the entire LAMP Corpus or more data might improve edit quality.
@@
7.3. What recommendations can we provide for future LLM-based writing support tools that aspire to improve co-writing experience?
@@
Eminent author Curtis Sittenfeld calls LLM writing the literary equivalent of fat-free cookies  [^11] https://www.nytimes.com/2024/08/20/opinion/beach-read-ai.html. LLMs are proficient at producing sentences that are grammatically correct and devoid of spelling errors. Beyond that, LLMs require extensive learning to effectively assist humans in improving their writing. In his essay Politics and English Language (Orwell, 2013), George Orwell said “Never use a long word when a short one will do.” LLM writing transgresses this simple rule by overusing lofty words. Clichés are bound to slip into even the best human writing, but when it comes to LLMs it simply cannot write without them. We believe this is partially a drawback of the technology behind LLMs. When a LLM calculates the probability of one word following another, clichés become very likely, because they’ve appeared so many times before. This explains why every other generated response is rife with clichés despite our prompt explicitly asking LLMs to avoid clichés and overused tropes (Table 13). LLMs need to learn how to identify and write without clichés such that it is engaging to every single reader. Overwriting is a bigger problem than underwriting. The rule for most writers is, “If in doubt, cut it.” (to Edit Your Own Writing, 2024) The Pulitzer Prize-winning writer John McPhee has called the process “writing by omission.” (McPhee, 2015). To become a better writer LLMs need to learn how to avoid unnecessary exposition. Last but not the least structure is what good writing hangs on (to Edit Your Own Writing, 2024). Long, run-on sentences are hard to read, and LLMs need to know when and how to split effectively to better manage flow and clarity.
@@ 8. Limitations
It should be noted that while paragraph-level editing provides a balance between granularity and context, it may miss broader structural or thematic issues that become apparent only when considering longer pieces of writing. Last but not least, our study relied on professional writers editing AI-generated text for monetary compensation, which may have influenced the quality and nature of the edits. Editing one’s own work typically involves more personal investment than editing text for pay, potentially leading to less motivation for substantial improvements (Byron and Khazanchi, 2012). Additionally, the repetitive nature of editing multiple AI-generated paragraphs could lead to fatigue, especially if the content is perceived as uninteresting or lacking in creativity. This fatigue could result in less thorough or thoughtful edits as the task progresses.
@@ 9. Conclusion
In this work, we present a comprehensive approach to mitigating idiosyncrasies and improving human-AI alignment in the writing process through expert editing. We i) develop a taxonomy of edit categories grounded in established writing practices, ii) create the LAMP corpus containing over 8,000 fine-grained edits by professional writers on LLM-generated text, and iii) design methods for automatic detection and rewriting of problematic spans. Our analysis reveals several key findings. Professional writers identify consistent categories of edits needed to improve AI writing. Surprisingly, there are no significant differences in perceived writing quality or types of edits needed across texts generated by different large language models (GPT-4, Claude 3.5, Llama 3.1). Automated methods using few-shot prompting are able to detect and rewrite problematic spans in LLM generated text, though far from matching human expert performance. Finally, in terms of preference evaluations, writers consistently rank text edited by other writers highest, followed by LLM-edited text, with unedited LLM-generated text ranking lowest. As AI text generation becomes more prevalent, developing robust editing and alignment techniques will be crucial to ensure AI systems produce high-quality writing that meets human standards and enhances creativity and linguistic diversity.
@@ A.1. Idiosyncracy Span Detection Prompt
{mdframed}
[ linecolor=black, linewidth=1pt, innertopmargin=6pt, innerbottommargin=6pt, innerrightmargin=10pt, innerleftmargin=10pt, frametitle=Idiosyncracy Span Detection Prompt, frametitlebackgroundcolor=black, frametitlefont=, ] You are given a paragraph of writing, and your goal is to provide feedback by selecting spans of text in the writing that could be improved, and assign each problematic span to an error category. Below, we list the 7 error categories that you can choose from.
@@
- ”Cliche”: The use of hackneyed phrases or overly common imagery that lack originality or depth.
@@
- ”Tense Consistency”: Comments pointing out inconsistencies in verb tense that need to be addressed for uniformity.
@@
- Number of Spans – You can provide feedback on multiple spans, and multiple spans can have the same category.
- Span must be verbatim – The span you select must be verbatim from the paragraph, otherwise, the feedback will not be provided to the user.
- No Overlap – Spans should not overlap, and one span should not include the other.
- Single Category – Each span should have exactly one category from the categories listed above.
@@ A.3. Rewriting Prompts
text after the edit, is its continuation. Simple way to ensure this is to make sure that the edited span has the same
@@
You will be given example of 25 paragraphs with text within <span></span>tags that shows poor sentence
@@
Your task will then be to suggest edits that rewrites the text within the span tags with better sentence structure
@@
PARAGRAPH READS COHERENTLY AND GRAMMATICALLY CORRECT. For instance if you replace

### v5

@@ 7.2. How well can LLMs mimic edits from writers?
Our preference ranking results in Section 6.3 indicate that automatically edited paragraphs frequently rank second and sometimes first. This raises questions about LLMs’ ability to analyze textual patterns and generate content closely resembling a given writer’s edit. For the span “Janet lay in bed each night, her mind a whirlpool of restless thoughts”, both LLM and writer identified it as cliché. The LLM edited it to “Janet lay in bed each night, unable to sleep” while the writer changed it to “Each night, Janet lay prone in her bed and unable to sleep”. LLMs can also split run-on sentences and improve poor structure. For “Sarah froze, realizing it was her high school sweetheart, Alex, whom she hadn’t seen in over a decade”, the LLM edited it to “Sarah froze. It was Alex, her high school sweetheart. She hadn’t seen him in over a decade”, similar to the writer’s edit. However, LLMs sometimes replace clichés with other clichés or fail to remove unnecessary exposition. The most challenging edit category is Lack of Specificity and Detail where LLMs often fail to add engaging details. For “Her irritation slowly morphed into a strange, disconnected calm”, the writer added “After all, the noise just meant that she wasn’t the only one awake at this hour.” The model’s edit was less effective: “Her irritation slowly morphed into a strange, disconnected calm. The repetitive thump-thump-thump became almost hypnotic, lulling her into a trance-like state”. Much evocative detail in human writing comes from lived sensory experiences. LLMs on the other hand lack the grounded understanding that helps humans select vivid, emotionally resonant details. Additionally, when we combine autoregressive objectives with post-hoc adjustment through RLHF ( typically designed to prevent toxic/harmful text generations) models often default to common, generic descriptions rather than specific ones since those are ”safer” predictions. However, we note one potential limitation in our experiments is our reliance on few-shot instructions, requiring the model to learn rewriting from only a few examples. Training on the entire LAMP Corpus or more data might improve edit quality.
@@
7.3. What recommendations can we provide for future LLM-based writing support tools that aspire to improve the co-writing experience?
@@
Eminent author Curtis Sittenfeld calls LLM writing the literary equivalent of fat-free cookies  [^12] https://www.nytimes.com/2024/08/20/opinion/beach-read-ai.html. LLMs are proficient at producing sentences that are grammatically correct and devoid of spelling errors. Beyond that, LLMs require extensive learning to effectively assist humans in improving their writing. In his essay Politics and English Language (Orwell, 2013), George Orwell said “Never use a long word when a short one will do.” LLM writing transgresses this simple rule by overusing lofty words. Clichés are bound to slip into even the best human writing, but when it comes to LLMs it simply cannot write without them. We believe this is partially a drawback of the technology behind LLMs. When an LLM calculates the probability of one word following another, clichés become very likely, because they’ve appeared so many times before. This explains why every other generated response is rife with clichés despite our prompt explicitly asking LLMs to avoid clichés and overused tropes (Table 14). LLMs need to learn how to identify and write without clichés such that it is engaging to every single reader. Overwriting is a bigger problem than underwriting. The rule for most writers is, “If in doubt, cut it.” (Guinness, 2024) The Pulitzer Prize-winning writer John McPhee has called the process “writing by omission.” (McPhee, 2015). To become a better writer LLMs need to learn how to avoid unnecessary exposition. Last but not least, structure is what good writing hangs on (Guinness, 2024). Long, run-on sentences are hard to read, and LLMs need to know when and how to split effectively to better manage flow and clarity.
@@
7.5. Why are there no significant differences in perceived writing quality or types of edits needed across texts generated by different large language models?
Figure 9. Preference Evaluation Results between GPT4o and GPT4-o1
The lack of significant quality differences between LLM writing is somewhat unexpected and warrants deeper examination. Several factors however may explain this surprising finding. Books play a crucial role in the training of generative AI systems. Their long, thematically consistent paragraphs provide information about how to generate coherent and fluent text. All three models (GPT-4, Claude 3.5, Llama 3.1) are pre-trained on Books3corpus (Kelly, 2023) which constitute a bigger portion of Pile (Gao et al., 2020) pre-training data. Additionally, all LLMs rely on Scale AI for collecting preference data and there is very likely a significant overlap in workers who annotate the preferences as recently discussed (Staff, 2024). Creative writing may also present an inherent ”quality ceiling” for current LLM architectures, where all models encounter similar limitations in maintaining a consistent narrative voice and avoiding clichés—a pattern also observed by (Tian et al., 2024) in their analysis of narrative generation. Our syntactic pattern analysis (Section 5.3) reveals remarkably similar templates and phrasings across models, further supporting the hypothesis of shared training foundations. While our methodology focused on concrete, actionable edit categories, we acknowledge that more subtle stylistic differences or subjective ”vibes” may exist between models not captured by our current taxonomy (Dunlap et al., 2024). The consistency in writing quality across models ultimately points to broader questions about the current limitations of LLMs in creative writing tasks, rather than differences between specific implementations. Recent preference evaluation results from OpenAI’s GPT4-o1 technical report (OpenAI, 2024) (See Figure 9) corroborate our findings where there are no significant differences in GPT4o vs GPT4o1 for personal writing or editing text.
@@ 8. Limitations
It should be noted that while paragraph-level editing provides a balance between granularity and context, it may miss broader structural or thematic issues that become apparent only when considering longer pieces of writing. Last but not least, our study relied on professional writers editing AI-generated text for monetary compensation, which may have influenced the quality and nature of the edits. Editing one’s own work typically involves more personal investment than editing text for pay, potentially leading to less motivation for substantial improvements (Byron and Khazanchi, 2012). Additionally, the repetitive nature of editing multiple AI-generated paragraphs could lead to fatigue, especially if the content is perceived as uninteresting or lacking in creativity. This fatigue could result in less thorough or thoughtful edits as the task progresses. Finally, another potential limitation of LLM-based editing is the risk of hallucinated and factually inconsistent information being introduced during the editing process. We did not study hallucinations as part of our work, due to the chosen domain focus of fictional writing which has less stringent factuality requirements, yet prior work has documented that LLM-based text editing can introduce factual errors (Laban et al., 2023a).
@@ 9. Conclusion
In this work, we present a comprehensive approach to mitigating idiosyncrasies and improving human-AI alignment in the writing process through expert editing. We i) develop a taxonomy of edit categories grounded in established writing practices, ii) create the LAMP corpus containing over 8,000 fine-grained edits by professional writers on LLM-generated text, and iii) design methods for automatic detection and rewriting of problematic spans. Our analysis reveals several key findings. Professional writers identify consistent categories of edits needed to improve AI writing. Surprisingly, there are no significant differences in perceived writing quality or types of edits needed across texts generated by different large language models (GPT-4, Claude 3.5, Llama 3.1). Automated methods using few-shot prompting can detect and rewrite problematic spans in LLM-generated text, though far from matching human expert performance. Finally, in terms of preference evaluations, writers consistently rank text edited by other writers highest, followed by LLM-edited text, with unedited LLM-generated text ranking lowest. As AI text generation becomes more prevalent, developing robust editing and alignment techniques will be crucial to ensure AI systems produce high-quality writing that meets human standards and enhances creativity and linguistic diversity.
@@ A.1. Idiosyncracy Span Detection Prompt
Idiosyncracy Span Detection Prompt
You are given a paragraph of writing, and your goal is to provide feedback by selecting spans of text in the writing that could be improved and assign each problematic span to an error category. Below, we list the 7 error categories that you can choose from.
@@
- ”Cliche”: The use of hackneyed phrases or overly common imagery that lacks originality or depth.
@@
- ”Tense Consistency”: Comments pointing out
inconsistencies in verb tense that need to be
addressed for uniformity.
@@
- Number of Spans – You can provide feedback on
multiple spans, and multiple spans can have the
same category.
- Span must be verbatim – The span you select must be
verbatim from the paragraph, otherwise, the feedback
will not be provided to the user.
- No Overlap – Spans should not overlap, and one
span should not include the other.
- Single Category – Each span should have exactly one
category from the categories listed above.
@@ A.3. Rewriting Prompts
text after the edit, is its continuation. A simple way to ensure this is to ensure that the edited span has the same
@@
You will be given examples of 25 paragraphs with text within <span></span>tags that shows poor sentence
@@
Your task will then be to suggest edits that rewrite the text within the span tags with better sentence structure
@@
PARAGRAPH READS COHERENTLY AND GRAMMATICALLY CORRECT. For instance, if you replace